# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis

### v1

@@ 5 Main results
In this proof we use the notation of left multiplication of functions by a matrix
$(Af)(u):=\inf\{f(x)\mid Ax=u\}.$
Note that $Af+\iota_{T_{a}}\equiv A(f+\iota_{A^{-1}(T_{a})})$. Indeed, both functions have the domain $(A\operatorname{dom}(f))\cap T_{a}$, where $A\operatorname{dom}(f):=\{Ax\mid x\in\operatorname{dom}(f)\}$, and $\forall y\in A\operatorname{dom}(f)\cap T_{a}$,
$(Af+\iota_{T_{a}})(y)$ $=\inf\{f(x)\mid Ax=y\}$
$=\inf\{f(x)+\iota_{Ax\in T_{a}}\mid Ax=y\}$
$=A(f+\iota_{A^{-1}(T_{a})})(y).$
Then, employing the change of variable $u=Ax$,
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)$ $=\inf_{u\in\mathbb{R}^{m}}(Af)(u)+g(u)$
$=\inf_{u\in\mathbb{R}^{m}}(Af+\iota_{T_{a}})(u)+(g+\iota_{T_{a}})(u)$
$=\inf_{u\in\mathbb{R}^{m}}(A(f+\iota_{A^{-1}{T_{a}}}))(u)+(g+\iota_{T_{a}})(u)$
$=\inf_{x\in\mathbb{R}^{n}}(f+\iota_{A^{-1}{T_{a}}})(x)+(g+\iota_{T_{a}})(Ax).$
Note that by Corollary 4.3, $\exists u\in\operatorname{relint}\operatorname{dom}(Af)\cap\operatorname{relint}\operatorname{dom}(g)$. Note that $\operatorname{dom}(Af)=A\operatorname{dom}(f)$, and by [14, Theorem 6.6], $\operatorname{relint}A\operatorname{dom}(f)=A\operatorname{relint}\operatorname{dom}(f)$. Therefore, $u\in A\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)$, i.e. there exists an $x\in\operatorname{relint}\operatorname{dom}(f)$ such that $Ax\in\operatorname{relint}\operatorname{dom}(g)$. With this, we satisfy a sufficient requirement of [14, Corollary 31.2.1], which guarantees strong duality and completes the proof. ∎
6 Nested normals and facial subspaces
Inspired by the lexicographic characterization of convex faces of [8], we develop a framework to describe the lattice of convex faces at a point $x\in C$. By using this theory, we will then be able to show the results of Section 4.
@@ 6 Nested normals and facial subspaces
We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
Theorem 6.1 (Facial subspace characterisation of convex faces).
@@
Let $C\subseteq X$. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
@@
This characterization, surprisingly, appears to be novel to the knowledge of the author. It allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set. We call such face-selecting subspaces affine facial subspaces.
Definition 6.1 (Affine facial subspace).
Let $C\subseteq X$ be convex. An affine subspace $L\subseteq V$ is an affine facial subspace of $C$ if $C\setminus L$ is convex.
The next result describes the relationship between affine facial subspaces and faces.
Corollary 6.1 (Facial subspace relation to faces).
@@
    For every affine facial subspace $L$ of $C$, the set $F:=C\cap L$ is a face of $C$.
@@
    Conversely, for every face $F$ of $C$, there exists an affine facial subspace $L$ such that $F=L\cap C$ (specifically, $L=\text{aff}(F)$).
@@
For the first part of the result, we show that if $L$ is an affine facial subspace, then $F:=C\cap L$ satisfies the conditions of Theorem 6.1.
@@
    which is convex by definition of $L$ as an affine facial subspace.
We show the second part of the result. For $F$ a face of $C$, by Theorem 6.1 $F=C\cap\operatorname{aff}(F)$ and $C\setminus F$ is convex. But then $C\setminus\operatorname{aff}(F)=C\setminus(\operatorname{aff}(F)\cap C)=C\setminus F$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, which makes $\operatorname{aff}(F)$ an affine facial subspace. ∎
In this work, we focus on “local” affine facial subspaces; the subset of affine facial subspaces that contain a point $x\in C$. These subsets have some additional nice algebraic properties. We find it useful to generalize this idea of “locality” to affine facial subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all affine subspaces that contain $S$. To emphasize this connection, we use the word “at” to describe the containment of $S$: ”a facial subspace is at SS”. The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
Any affine facial subspace $L$ at $S$ has its affine offset determined by $S$. Recall that LLL-L is simply the affine subspace $L$ translated in such a way that it becomes a (non-affine) subspace. With this, it becomes clear that L=(LL)+SL=(L-L)+S. We can always deduce the correct affine offset from in $S$. Therefore, we find it convenient to describe the set of affine facial subspaces that contain $S$ through their translated, non-affine counterparts.
Definition 6.2 (Facial subspace at S).
Let $C\subseteq X$ be a convex set and $S\subseteq X$ a non-empty subset. A subspace $U\subseteq X$ is defined to be a facial subspace of $C$ at $S$ if:
@@
    $U+S$ is an affine facial subspace of $C$.
The set of all facial subspaces of $C$ at $S$ is denoted $\mathcal{F}(C;S)$.
When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied. When $S$ is not a singleton, this assumption guarantees that $\forall s_{1},s_{2}\in S,U+s_{1}=U+s_{2}$. This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$. In other words, the affine facial subspace $U+S$ contains $S$.
@@
We can see that a normal cone is a dual object which “describes” the original set only up to its closure.
We require a dual object which has sufficient descriptive power to go beyond the closure, encoding information about the boundary. To this end, we introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
@@
We would also like for our normal cones to located at a set $S\subseteq C$, similarly to facial subspaces.
@@
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [7], where normal cones are at faces of polytopes, instead at points. The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at any point in the relative interior of a convex set $S\subseteq C$ will match the normal cone of $S$.
@@
Lemma 6.1 (nested normals in the relative interior are representative).
@@
That $N_{C}^{U}(S)\subseteq N_{C}^{U}(x)$ follows from simple inclusion $x\subseteq S$.
@@
Lemma 6.2 (Matching sets means paired normals).
@@
The tangent cone of a convex set $C$ at $x\in C$ is the smallest closed cone $T$ such that $C\subseteq x+T$. We can think of this cone as a local description of $C$ around $x$. The normal cone is the polar of hthe tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone. In contrast, nested normals in facial subspaces of $C$ at $x$ describe the cone of feasible directions.
@@
For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be $D_{C}(x):=\{v\in\mathbb{R}^{n}\mid\exists\lambda>0:\lambda v\in C\}$.
The closure of the cone of feasible directions is the tangent cone. The difference is that the cone of feasible directions captures additional information about the boundary of $C$ around $x$. The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in facial subspaces.
@@
$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
@@
Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$. By the definition of conditional normals, $T\subseteq T^{\prime}$. For the containment ($\supseteq$), we argue by contradiction: we assume that there is some $x\in T^{\prime}\setminus T$. If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [14, Theorem 11.3]). Any such hyperplane induces a nested normal of the form $v\in N_{T}^{\mathbb{R}^{n}}(0)$, for which $v^{\perp}$ is the hyperplane, and $\langle v,x\rangle>0$. But then $x\notin T^{\prime}$, and so $x\notin\mathrm{int}(T^{c})$. Since we assumed that $x\notin T$, the only possibility is that $x\in\partial T$. Then $x$ is contained in a supporting hyperplane, so there is some nested normal $v\in N_{T}^{\mathbb{R}^{n}}(0)$ such that $\langle v,x\rangle=0$. Let $T_{1}=T\cap v^{\perp}$, and narrow the ambient space to the subspace $T_{1}$. Inside $T_{1}$, we repeat the same argument, and keep narrowing the ambient space in this manner. This process must eventually reduce the ambient space to $\operatorname{\mathrm{span}}(x)$, because $\mathbb{R}^{n}$ is finite-dimensional. In this one-dimensional space, with the narrowed cone $T_{\ell}=T\cap\operatorname{\mathrm{span}}(x)$, note that $T_{\ell}$ is a closed cone. Therefore, either $x\in T_{\ell}$ or $x\in\mathrm{int}(T_{\ell}^{c})$, both of which yields a contradiction, from which the result follows. ∎
Another important property of nested normals is that they describe the containment structure of facial subspaces, which is closely related to the lattice structure of the faces. We make this precise in the next two results.
Lemma 6.3 (Nested normals induce facial subspaces).
@@
Lemma 6.4 (Nested normals induce facial subspaces).
@@
We make two observations.
First, since $v\in N_{C}^{U}(S)$ and $V\subseteq U$, $\forall y\in C\cap(V+S)$, $\forall s\in S,$ we have that $\langle y,v\rangle\leq\langle s,v\rangle$. Second, this inequality is tight precisely for those values of $y$ that are also contained in $v^{\perp}+S$, i.e.,
$\{y\in C\cap(V+S)\mid\forall s\in S,\langle y,v\rangle=\langle s,v\rangle\}$ (5)
$=$ $((V+S)\cap C)\cap(v^{\perp}+S)=(V\cap v^{\perp}+S)\cap C.$ (6)
@@
The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, i.e., $V\cap v^{\perp}\in\mathcal{F}(C;S)$.
We next show that all facial subspaces are “reached” by nested normals in facial subspaces.
Lemma 6.5 (Completeness of conditional normal proof composition).
@@
This result is similar to the lexicographic characterization of the faces of convex sets in [8, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face. Let us state a technical lemma that we will need for the proof.
Lemma 6.6 (Affine slices preserve facial subspaces).

### v3

@@ 5.3 Convex optimization
By Corollary 4.1, $(A\operatorname{dom}(f))\cap\operatorname{dom}(g)\subseteq T_{a}$, and therefore $\iota_{T_{a}}(Ax)=0$ for all $x\in\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)$. Additionally,
$\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)\subseteq A^{-1}(A\operatorname{dom}(f)\cap\operatorname{dom}(g))\subseteq A^{-1}T_{a},$
which means that $\iota_{A^{-1}T_{a}}(x)$ is 0 on $\operatorname{dom}(f+g\circ A)$. And for any $x\notin\operatorname{dom}(f+g\circ A)$, it is easily checked that both sides of Equation 6 evaluate to $\infty$, so Equation 6 holds for all $x\in\mathbb{R}^{n}$.
For the second part of the result,
$A\operatorname{relint}\operatorname{dom}(f+\iota_{A^{-1}(T_{a})})\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))$ (7)
$=$ $\operatorname{relint}(A\operatorname{dom}(f+\iota_{A^{-1}(T_{a})}))\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))$ (8)
$=$ $\operatorname{relint}(A(\operatorname{dom}(f)\cap A^{-1}(T_{a})))\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})$ (9)
$=$ $\operatorname{relint}((A\operatorname{dom}(f))\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (10)
where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and the last inequality is guaranteed by Corollary 4.5. ∎
We present an exact Fenchel–Rockafellar dual for 5.1: a dual problem for which strong duality always holds and the dual attained.
Corollary 5.5 (Exact Fenchel-Rockafellar dual).
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ l.s.c. proper convex functions, and a matrix $A\in\mathbb{R}^{n\times m}$, let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
$\inf_{x\in\mathbb{R}^{n}}[f(x)+g(Ax)]=-\min_{y\in\mathbb{R}^{m}}[(f+\iota_{A^{-1}(T_{a})})^{*}(A^{T}y)+(g+\iota_{T_{a}})^{*}(-y)].$
Proof of Corollary 5.5.
The result follows from Theorem 5.1 with the use of [rockafellarConvexAnalysis1970, Corollary 31.2.1]. ∎
Definition 5.1 (Minimizer).
A minimizer of a given convex program is any point $x\in\mathbb{R}^{n}$ that attains the infimum of the program, provided that the infimum is finite.
Next, we provide a characterization of the minimizers of 5.1, i.e., KKT optimality conditions. Our KKT condition holds without any further assumptions on the program; we do not require strong consistency.
Corollary 5.6 (KKT conditions).
Consider 5.1, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.1 if, and only if, the following conditions hold:
  1. 1.
    $Ax\in\partial(g+\iota_{T_{a}})^{*}(y).$
  2. 2.
    $A^{T}y\in\partial(f+\iota_{A^{-1}(T_{a})})(x)$.
Proof of Corollary 5.6.
From Theorem 5.1, the result follows from application of [rockafellarConvexAnalysis1970, Corollary 31.3.1] to the functions $f+\iota_{A^{-1}(T_{a})}$ and $g+\iota_{T_{a}}$. ∎
Infeasible case as convention
In the above results, the domains of the function under consideration may not intersect, in which case $f+g\circ A\equiv\infty$. Intersection is required for the existence of a meaningful joint supporting subspace, so it might be reasonable to make this non-trivial intersection a requirement. This kind of requirement was dubbed a “universal qualification condition” in the context of conic programming [tuncelStrongDualityMinimal2012]. We observe that such a condition is not a true limitation of the theory, because the behaviour outside this condition is known and simple. Therefore, we instead use a convention: whenever $C\cap D=\emptyset$, we let $T_{a}(C,D):=\emptyset$ as well. As a result, $f+\iota_{A^{-1}(T_{a})}\equiv\infty$, and the equality in Corollary 5.5 trivially holds because both sides evaluate to $\infty$, with the value of $\infty$ attained in the dual.
5.3.1 Facial reduction
We derive two known facial reduction results, thereby showing that the joint supporting subspace is indeed a generalization of facial reduction.
Problem 5.2 (Conic program).
For $K\subseteq\mathbb{R}^{n}$ a closed convex cone, $A\in\mathbb{R}^{m\times n}$ a matrix, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$,
$\mathop{\mathrm{minimize}}_{x\in\mathbb{R}^{n}}\langle x,c\rangle\quad\mathrm{s.t.}\quad Ax=b,\quad x\in K.$
We apply Corollary 5.5 and Corollary 5.6 to this problem, yielding analogues of [tuncelStrongDualityMinimal2012, Theorem 4.3] and [tuncelStrongDualityMinimal2012, Corollary 4.4]. The derived results differ from those in [tuncelStrongDualityMinimal2012] only in that our primal problem is their dual.
Corollary 5.7 (Exact conic dual).
For $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, $c\in\mathbb{R}^{n}$, and a closed convex cone $K\subseteq\mathbb{R}^{n}$,
$\inf\{\langle x,c\rangle\mid x\in\mathbb{R}^{n},Ax=b,\text{ and }x\in K\}$
$\max\{\langle b,y\rangle\mid y\in\mathbb{R}^{m}\text{ and }A^{T}y-c\in F^{\circ}\},$
where $F=K\cap A^{-1}(H_{AK}(\{b\}))$.
Proof of Corollary 5.7.
We use Corollary 5.5 with $f(x):=\langle c,x\rangle+\iota_{K}(x)$, and $g:=\iota_{\{b\}}$. We find
$T:=T(A\operatorname{dom}(f),\operatorname{dom}(g))=T(AK,\{b\})=H_{AK}(\{b\}),$
where the last equality follows from Theorem 4.1. Since $AK$ is a cone, any face includes the origin, and therefore $0\in H_{AK}(\{b\})=T_{a}$, which means that $T_{a}=T$. The functions in the dual program are
$f(x)+\iota_{A^{-1}(T)}(x)$ $=\langle c,x\rangle+\iota_{K}(x)+\iota_{A^{-1}(H_{AK}(\{b\}))}(x)$
$=\langle c,x\rangle+\iota_{K\cap A^{-1}(H_{AK}(\{b\}))}(x)$
and
$g(Ax)+\iota_{T}(Ax)=\iota_{\{b\}}(Ax)+\iota_{H_{AK}(\{b\})}(Ax)=\iota_{\{b\}}(Ax).$
Let $F=K\cap A^{-1}(H_{AK}(\{b\}))$, the subset of $K$ selected by the joint supporting subspace. The set $F$ is a cone, because $H_{AK}(\{b\})\cap AK$ is a face of the cone $AK$ by Definition 6.5 and Corollary 6.1, and $F$ is therefore itself a cone.
Then with [rockafellarConvexAnalysis1970, Theorem 12.3, Theorem 14.1] we compute that
$(f+\iota_{A^{-1}(T)})^{*}(A^{T}y)+(g+\iota_{T})^{*}(-y)$ $=(\langle c,\cdot\rangle+\iota_{F})^{*}(A^{T}y)+(\iota_{\{b\}})^{*}(-y)$
$=\iota_{F^{\circ}}(A^{T}y-c)-\langle b,y\rangle.$
We thereby find the dual problem, for which, by Corollary 5.5, strong duality holds, and the dual is attained. ∎
Corollary 5.8 (Conic KKT conditions).
A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.2 if, and only if, there exists some $y\in\mathbb{R}^{m}$ such that all of the following hold.
  1. 1.
    $Ax=b$.
  2. 2.
    $A^{T}y-c\in F^{\circ}$.
  3. 3.
    $\langle x,A^{T}y-c\rangle=0$.
Proof of Corollary 5.8.
We apply Corollary 5.6 to the case where $f(x):=\langle c,x\rangle+\iota_{K}(x)$ and $g:=\iota_{\{b\}}$. As in the proof of Corollary 5.7, we find that $f(x)+\iota_{A^{-1}(T_{a})}=\langle c,x\rangle+\iota_{F}(x)$ and $g(\bar{y})+\iota_{T_{a}}(\bar{y})=\iota_{\{b\}}(\bar{y})$.
Then $(g+\iota_{T_{a}})^{*}(y)=\langle b,y\rangle$, and the first KKT condition, that $\partial(g+\iota_{T_{a}})^{*}(y)=\{b\}$ reduces to $Ax=b$. The second KKT condition is
$A^{T}y$ $\in\partial(f+\iota_{A^{-1}(T_{a})})(x)$
$=\partial(\langle c,\cdot\rangle+\iota_{F})(x)=c+N_{F}(x).$
But since $N_{F}(x)=F^{\circ}\cap x^{\perp}$, this second condition holds if, and only if, both $x\perp A^{T}y-c$ and $A^{T}y-c\in F^{\circ}$. ∎
6 A framework for local face lattices
@@
Inspired by the lexicographic characterization of convex faces of [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_], we develop a framework describing the lattice of convex faces at a point $x\in C$, which we will then use to show the results of Section 4. We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
Theorem 6.1 (Characterization of convex faces by supporting subspaces).
@@
Let $C\subseteq X$ be a convex set. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
@@
Theorem 6.1 motivates the next mathematical object.
Definition 6.1 (Supporting subspace).
Let $C\subseteq X$ be convex. A supporting subspace $L\subseteq X$ of $C$ is an affine subspace such that $C\setminus L$ is convex.
Though supporting subspaces are affine subspaces, we omit the adjective “affine” for brevity, mirroring the term “supporting hyperplane”. The notion of supporting subspaces seems to be surprisingly sparse in the literature; the only previous appearance of this notion the author could find is in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4]. This idea is of central importance to this paper; it allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set. The next result describes the relationship between supporting subspaces and faces.
Corollary 6.1 (Relation of supporting subspaces to faces).
@@
    For any affine subspace $L\subseteq X$,
[…]
@@
    For any subset $F\subseteq C$,
[…]
@@
1), $(\implies)$ we show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
@@
    which is convex by definition of $L$ as a supporting subspace.
[…]
2), ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace. Indeed, by Theorem 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex. Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
[…]
In this work, we focus on “local” supporting subspaces; the subset of supporting subspaces that contain a point $x\in C$. These subsets have some additional nice algebraic properties. We find it useful to generalize this idea of “locality” to supporting subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all supporting subspaces of $C$ that also contain $S$. To emphasize this connection, we say that a supporting subspace is ”at SS” to describe its containment of $S$. The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
Any supporting subspace $L$ at $S$ has its affine offset determined by $S$. Recall that LLL-L is the tangent space of $L$, i.e., the translation of $L$ that becomes a (non-affine) subspace. With this, it becomes clear that L=(LL)+SL=(L-L)+S. We can always deduce the correct affine offset from $S$ alone. Therefore, we find it convenient to describe the set of supporting subspaces that contain $S$ through their tangent spaces. We henceforth call the tangent spaces of supporting subspaces “supporting subspaces” as well, and whether we mean the affine subspace or its tangent space will be clear from context.
Definition 6.2 (Supporting subspace at S).
Let $C\subseteq X$ be a convex set and $S\subseteq X$ a non-empty subset. We call a subspace $U\subseteq X$ a supporting subspace of $C$ at $S$ if:
@@
    $U+S$ is an (affine) supporting subspace of $C$ as defined in Definition 6.1.
We define $\mathcal{F}(C;S)$ to be the set of all subspaces $U$ satisfying both conditions above for fixed $C$ and $S$.
When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied. When $S$ is not a singleton, this assumption guarantees that $\forall s_{1},s_{2}\in S,U+s_{1}=U+s_{2}$. This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$. In other words, the supporting subspace $U+S$ contains $S$.
@@
We can see that a normal cone is a dual object which “describes” the original cone only up to its closure.
We require a dual object which has sufficient descriptive power to go beyond the closure, by encoding information about the boundary. To this end, we introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
@@
We would also like for our normal cones to located at a set $S\subseteq C$, similarly to supporting subspaces.
@@
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are at faces of polytopes, instead at points. The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at any point in the relative interior of a convex set $S\subseteq C$ will match the normal cone of $S$.
@@
Lemma 6.1 (Normal cones in the relative interior are representative).
@@
That $N_{C}^{U}(S)\subseteq N_{C}^{U}(x)$ follows from simple inclusion $x\in S$.
@@
Lemma 6.2 (Normal cone everywhere is the orthogonal subspace).
@@
The tangent cone of a convex set $C$ at $x\in C$ is the smallest closed cone $T$ such that $C\subseteq x+T$. We can think of this cone as a local description of $C$ around $x$. The normal cone is the polar of the tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone. In contrast, nested normals in supporting subspaces of $C$ at $x$ describe the cone of feasible directions.
@@
For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be
$D_{C}(x):=\operatorname{cone}(C-x).$
The closure of the cone of feasible directions is the tangent cone. The difference is that the cone of feasible directions captures additional information about the boundary of $C$ around $x$. The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in supporting subspaces.
@@
$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
@@
Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$. By the definition of nested normals, $T\subseteq T^{\prime}$. For the containment ($\supseteq$), we argue by contradiction: we assume that there is some $x\in T^{\prime}\setminus T$. If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [rockafellarConvexAnalysis1970, Theorem 11.3]). Any such hyperplane induces a nested normal of the form $v\in N_{T}^{\mathbb{R}^{n}}(0)$, for which $v^{\perp}$ is the hyperplane, and $\langle v,x\rangle>0$. But then $x\notin T^{\prime}$, and so $x\notin\mathrm{int}(T^{c})$. Since we assumed that $x\notin T$, the only possibility is that $x\in\partial T$. Then $x$ is contained in a supporting hyperplane, so there is some nested normal $v\in N_{T}^{\mathbb{R}^{n}}(0)$ such that $\langle v,x\rangle=0$. Let $T_{1}=T\cap v^{\perp}$, and narrow the ambient space to the subspace $T_{1}$. Inside $T_{1}$, we repeat the same argument, and keep narrowing the ambient space in this manner. This process must eventually reduce the ambient space to $\operatorname{\mathrm{span}}(x)$, because $\mathbb{R}^{n}$ is finite-dimensional. In this one-dimensional space, with the narrowed cone $T_{\ell}=T\cap\operatorname{\mathrm{span}}(x)$, note that $T_{\ell}$ is a closed cone. Therefore, either $x\in T_{\ell}$ or $x\in\mathrm{int}(T_{\ell}^{c})$, both of which yields a contradiction, from which the result follows. ∎
Another important property of nested normals is that they describe the containment structure of supporting subspaces, which is closely related to the lattice structure of the faces. We make this precise in the next two results.
Theorem 6.3 (Nested normals induce supporting subspaces).
@@
Proof of Theorem 6.3.
@@
We make two observations. First, since $v\in N_{C}^{U}(S)$ and $V\subseteq U$, $\forall y\in C\cap(V+S)$, $\forall s\in S,$ we have that $\langle y,v\rangle\leq\langle s,v\rangle$. Second, this inequality is tight precisely for those values of $y$ that are also contained in $v^{\perp}+S$, i.e.,
$\{y\in C\cap(V+S)\mid\forall s\in S,\langle y,v\rangle=\langle s,v\rangle\}$
$=$ $((V+S)\cap C)\cap(v^{\perp}+S)=(V\cap v^{\perp}+S)\cap C.$
@@
The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, which by Corollary 6.1 implies that $V\cap v^{\perp}\in\mathcal{F}(C;S)$. ∎
We next show that all supporting subspaces are “reached” by nested normals in supporting subspaces.
Theorem 6.4 (Completeness of the composition of nested normals).
@@
This result is similar to the lexicographic characterization of the faces of convex sets in [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face. Let us state a technical lemma that we will need for the proof.
Lemma 6.3 (Affine slices preserve supporting subspaces).